OpenStudy (anonymous):

Simplify -11(Sqrt)12

OpenStudy (anonymous):

\[-11\sqrt{12}\]

OpenStudy (anonymous):

112* sorry not 12

jhonyy9 (jhonyy9):

how is right ?

jhonyy9 (jhonyy9):

sure

jhonyy9 (jhonyy9):

-sqrt112 ???

OpenStudy (anonymous):

\[-11\sqrt{112}\]

jhonyy9 (jhonyy9):

-11sqrt(2*56) = -11sqrt(2*8*7) = -11sqrt(16*7) = -11*4sqrt7 = -44sqrt7